OpenStudy (anonymous):

identify the domain and range and sketch the graph of y=1/ x-2

OpenStudy (anonymous):

\[y=\frac{1}{x-2}\]?

OpenStudy (anonymous):

yup !

OpenStudy (anonymous):

to find the domain, solve \[x-2=0\] in your head and then say "all real number except \(x=2\)"

OpenStudy (anonymous):

as for the range, a fraction is only zero if the numerator is your numerator is 1, so it is never zero range: all real numbers except 0

OpenStudy (anonymous):

and how would we write those out?

OpenStudy (anonymous):

i just wrote them out what would you like to write?

OpenStudy (anonymous):

I meant like putting it in parenthesis

OpenStudy (anonymous):

oh like \[(-\infty,2)\cup(2, \infty)\]?

OpenStudy (anonymous):

Oh!! totally got it now! Thanks!! :)

OpenStudy (anonymous):

that is a kind of long winded way of saying \(x\neq 2\)

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

one more question, how would we go about graphing that ?

OpenStudy (anonymous):

@satellite73